OpenStudy (anonymous):

Determine the values of C that makes the following function continuous F(x)=3x+c, when x is less than or equal to 2 F(x)=C^2-2x+4, when x is greater than 2 ..I got it down to 6-c=c^2 but I don't know how to simplify that.

OpenStudy (anonymous):

Skip the first part (until you see red) if you don't want to read it; I'm just recreating what your work probably looks like. \[f(x)=\begin{cases}3x+c&\text{if $x\le2$}\\ c^2-2x+4&\text{if $x>2$}\end{cases}\] For f to be continuous at x = 2, the following must be satisfied: \[\lim_{x\to2^+}f(x)=\lim_{x\to2^-}f(x)\] The one-sided limits are as follows: \[\begin{align*}\lim_{x\to2^+}f(x)&=\lim_{x\to2^+}\left(c^2-2x+4\right)\\ &=c^2-4+4\\ &=c^2\end{align*}\] and \[\begin{align*}\lim_{x\to2^-}f(x)&=\lim_{x\to2^-}\left(3x+c\right)\\ &=\color{red}{6+c}\\\end{align*}\] So you made a small mistake when you evaluated the second limit. Now, setting the limits equal to each other, \[c^2=6+c\\ c^2-c-6=0\] Have you tried the quadratic formula?